OpenStudy (anonymous):

how do you solve an equation if it's wrote like this.. 5x-3=4(x+2) ? if it has the (x+2) at the end?

OpenStudy (anonymous):

You need to expand 4(x+2) to get rid of parentheses. So, you get 4x + 4*2 = 4x+8. Then, solve the rest.

OpenStudy (anonymous):

distribute the 4

OpenStudy (anonymous):

5x-3=4x+8 x-3=8 x=11

OpenStudy (anonymous):

the plug back in the value of x

OpenStudy (anonymous):

whis is 52=52